# TaoSR-AGRL: Adaptive Guided Reinforcement Learning Framework for E-commerce Search Relevance

### v2

@@ 3.3. Adaptive Guided Replay
While Rule-aware Reward Shaping provides fine-grained feedback, reward sparsity remains a significant challenge, particularly for novel, long-tail queries involving composite rules. This issue is best illustrated by the asymmetry in e-commerce search rules: a “cashmere clothing” is rated 4-Excellent only if its cashmere content exceeds 50%, whereas for “chiffon”, mere presence suffices. Such composite, asymmetric constraints demand multi-step, fine-grained reasoning, yet models refined via SFT or DPO typically exhibit contracted action spaces and limited exploration. As a result, when an LLM happens to output the correct final label without articulating the intermediate logic, it receives no credit for the critical sub-tasks. The resulting learning signal is sparse precisely where reasoning is most demanding, degrading credit assignment and hindering convergence on complex e-commerce rules.
@@
To prevent the model from engaging in reward hacking or degrading its reasoning abilities when given complete answers, we adopt a “minimal sufficient guidance” strategy. AGR operationalizes this by performing a fine-grained, on-demand diagnosis. Leveraging the multi-dimensional signals from our rule-aware reward, AGR first calculates the in-batch average accuracy for each reasoning dimension (e.g., category, attribute, and relevance) and compares it against a guidance trigger threshold $\tau$. For any dimension where the accuracy falls below this threshold, AGR generates targeted guidance $\mathcal{G}(x,y)$ for the selected sparse-reward samples $(x,y)$. This guidance is then formatted into a text prompt and appended to the original input $P(x)$ to form an augmented prompt $P^{\prime}(x)$. The model performs a “guided replay” by regenerating its output based on $P^{\prime}(x)$, allowing it to explore higher-reward reasoning trajectories and fill gaps in its understanding of fine-grained discriminative logic. This dimension-specific intervention effectively mitigates reward sparsity while preserving the model’s autonomous reasoning process and suppressing reward hacking.
4. Experiment
@@ 4.1. Experimental Setup
For offline evaluation, we report per-class F1 for Class-1 (1-Irrelevant), Class-2 (2-Mismatch), Class-3 (3-Related), and Class-4 (4-Excellent), along with Good F1, Macro F1, and Accuracy. Good F1 is computed by collapsing Class-3 and Class-4 into a single “Good” category, aligning our evaluation protocol with the online business objective where relevance levels 3-Related and 4-Excellent are treated as a unified “Good Tie”.
To validate the real-world impact of our model beyond offline metrics, we conducted online side-by-side human evaluations. This rigorous process is designed to capture user-facing performance differences by directly assessing model outputs on three principal metrics:
  • Good/Same/Bad (GSB): This metric quantifies the relative performance gain of our model via a pairwise comparison. For the same query, human annotators are presented with two Search Engine Result Pages (SERPs) side-by-side—one from our model and one from the baseline—and label our model’s result as “Good” (superior), “Same” (equivalent), or “Bad” (inferior). A GSB score of +xx% signifies that our model produced a superior SERP for xx% more queries compared to the baseline.
  • Query Goodrate: This is a holistic, SERP-level metric assessing the absolute quality of search results. Annotators classify the overall utility of an entire SERP into “Good”, “Mid”, or “Bad” tiers. The final metric is the percentage of queries whose SERPs are judged as either “Good” or “Mid”, providing a high-level measure of user satisfaction per query.
@@
All experiments related to TaoSR-AGRL were implemented based on the open-source reinforcement learning framework ROLL(wang2025reinforcementlearningoptimizationlargescale). We used a training batch size of 64, a maximum input length of 2048, and a maximum generation length of 4096. The learning rate was set to 1e-6, with 16 gradient accumulation steps. For each input sample, 16 candidate responses were generated using a sampling temperature of 0.99 and a top-k of 100. For difficulty filtering, an online difficulty threshold range of [0.01, 0.9] was used. To enhance training stability, the PPO clipping range was set to $1\pm 0.2$, value function clipping to 0.5, and Advantage value clipping to $\pm 2.0$. The guidance trigger threshold $\tau$ was 0.1, and the hyperparameter weights $w_{\text{cate}}$, $w_{\text{attr}}$, and $w_{\text{reasoning}}$ were set to 0.4, 0.4, and 0.2 through preliminary experiments to balance different reward components effectively.
@@ 4.4. Ablation Study
We conduct a series of ablation studies on the W-Eval benchmark to systematically analyze the performance contributions of each component in TaoSR-AGRL. The main results are summarized in Table 2. In our analysis, the suffix -RRS denotes the variant equipped with Rule-aware Reward Shaping. The suffix -FG (Fixed Guidance) denotes a special case of our Adaptive Guided Replay where the guidance is fixed to the ground-truth relevance labels.
@@
In reinforcement learning, policy entropy is a critical measure of exploration. High entropy is essential for deep reasoning tasks, as it encourages the model to explore a diverse set of inferential strategies. We compare TaoSR-AGRL against TaoSR-AGRL-Static, a variant that applies guidance indiscriminately to all samples. As illustrated in Figure 2(c), TaoSR-AGRL consistently maintains a higher policy entropy. This suggests that excessive or non-adaptive guidance can lead to policy collapse(juliani2024studyplasticitylossonpolicy): the model becomes overly reliant on the guidance, causing its output distribution to become over-confident and collapse into a narrow set of familiar responses. This dependency ultimately stifles the model’s exploratory capabilities and impairs its ability to generalize.
Table 3. Effect of $\tau$ on the In-the-Wild evaluation set.
Threshold ($\tau$) Macro F1 Accuracy
0 60.23 73.58
0.1 61.53 74.28
0.2 59.91 74.10
0.3 59.40 73.77
0.4 59.45 73.64
0.5 58.70 74.27
@@
To determine the optimal frequency of intervention, we analyze the model’s overall performance across different values of the replay trigger threshold $\tau$. As presented in Table 3, performance peaks at $\tau=0.1$. This reveals a critical trade-off. A low $\tau$ provides insufficient correction for challenging instances, hindering learning. Conversely, a high $\tau$ leads to over-reliance on external guidance, risking the policy collapse phenomena discussed previously. This over-reliance can impair the model’s intrinsic reasoning ability, even on simpler tasks.
Synergy of Gating and Reasoning Rewards.
To dissect our Rule-aware Reward Shaping, we ablated its two pillars: the validity gate and the reasoning reward ($R_{\mathrm{reason}}$). Beyond end-task metrics, we use the Rule Adherence Rate (RAR) to measure if the model’s reasoning is faithful to its own intermediate steps. Ablating the hard validity gate—either by removing or softening it—introduces reward leakage into invalid trajectories. This not only degrades end-task performance but also lowers RAR, thereby confirming that a strict gate is essential for providing a clean, focused training signal. More critically, removing $R_{\mathrm{reason}}$ causes a steep decline in RAR while leaving end-task metrics largely intact. This result provides clear evidence of reward hacking: the model learns to find the correct answer via logically flawed shortcuts. This demonstrates the synergy between our components: the hard gate ensures learning is confined to successful outcomes, while $R_{\mathrm{reason}}$ ensures the reasoning process leading to those outcomes is logically sound and faithful. A detailed breakdown of these ablations is available in Appendix A.
@@ 4.5. Online Evaluation
To assess the real-world effectiveness of the proposed model, we conducted large-scale online side-by-side human evaluations. TaoSR1 was used as the baseline bucket, whereas TaoSR‑AGRL was employed as the test bucket. For 2,000 authentic user queries, the top‑10 retrieval results of both systems were compared item‑by‑item. The results indicated that TaoSR‑AGRL consistently outperformed the baseline bucket, achieving average improvements of 7.11% in GSB, 1.74 pt in Query Goodrate and 3.74 pt in Item Goodrate. Gains were particularly pronounced on reasoning‑intensive tasks (e.g., Q&A and knowledge‑based queries), indicating that the relevance model effectively integrates domain‑specific e‑commerce knowledge with rule‑driven reasoning to handle complex cases.
However, initial online deployment revealed a slight decline in critical business metrics—including order volume and GMV—despite the observed relevance gains. Diagnostic analysis traced the issue to the upstream recall stage, which disproportionately favored items with high semantic relevance but low sales velocity, thereby impairing conversion efficiency. This phenomenon revealed a classic trade-off: a system optimized solely for semantic relevance can inadvertently penalize items with strong commercial appeal but a less-perfect textual match. To address this, we introduced a multi-path recall mechanism incorporating personalized signals, alongside a pre-ranking stage that jointly optimizes for relevance and commercial potential. These enhancements restored business metrics to statistical parity with the baseline while preserving the documented relevance improvement. The results highlight the necessity of jointly designing retrieval and ranking to ensure semantic quality and commercial objectives are simultaneously satisfied in end-to-end search systems.
@@ 5. Conclusion
In this paper, we proposed TaoSR-AGRL, an adaptive guided reinforcement learning framework that systematically addresses the critical challenges of reward sparsity and exploration stagnation in LLM-based e-commerce search relevance. Our approach integrates two key innovations: Rule-aware Reward Shaping, which decomposes sparse terminal outcomes into dense, multi-dimensional rewards to enforce faithful reasoning, and Adaptive Guided Replay, which provides targeted, on-demand guidance to help the policy explore high-value trajectories for challenging long-tail instances. Extensive offline experiments and online side-by-side human evaluations demonstrated that TaoSR-AGRL significantly outperforms strong baselines, improving both relevance accuracy and rule adherence while ensuring stable training. The resulting model has been successfully deployed in Taobao Search, serving hundreds of millions of users and providing a practical blueprint for applying reinforcement learning to build robust, interpretable, and industrial-scale reasoning systems.
@@ Appendix A Ablation Study on Rule-aware Reward Shaping
We conducted an ablation study to validate the design of our Rule-aware Reward Shaping. The study dissects two key components: (1) the validity gate ($\mathbb{I}_{\mathrm{Gate}}$), which selectively applies rewards, and (2) the reasoning reward ($R_{\mathrm{reason}}$), which promotes logical consistency. TaoSR-AGRL uses a hard gate (Gating Coefficient $\lambda=0$), meaning no reward is given if the trajectory fails to satisfy the gating condition. To assess whether the model derives relevance in accordance with the operational rules, we report the Rule Adherence Rate (RAR). RAR measures the proportion of predictions whose final Relevance tier is consistent with the predicted Category and Attribute tiers under the Relevance Derivation Rules in Table 6. As such, RAR captures rule-following behavior and is complementary to end-task metrics (F1, Accuracy): a model may achieve high Accuracy while violating the rules internally, which RAR is designed to reveal.
@@
As shown in Table 5, TaoSR-AGRL outperforms all ablated variants and attains the highest RAR. We observe three consistent trends: (1) relaxing or removing the gate causes reward leakage into invalid trajectories, which not only degrades task metrics but also reduces RAR, indicating poorer rule adherence; (2) increasing the softness of the gate exacerbates this leakage and further erodes internal consistency, underscoring the necessity of a hard gate to concentrate learning on valid reasoning chains; (3) removing $R_{\mathrm{reason}}$ yields notably lower RAR despite similar headline metrics, revealing shortcut solutions—a form of reward hacking—that bypass the rules. Together, these findings show that the hard gate provides a clean supervision boundary, while $R_{\mathrm{reason}}$ regularizes the solution space to faithful, rule-consistent derivations, improving both utility and reliability.
@@
Appendix B Relevance Derivation Rules
Table 6 summarizes the Relevance Derivation Rules currently employed in Taobao. Given specific Category and Attribute tiers, the corresponding relevance tier can be determined directly from the table.
@@ Appendix B Relevance Derivation Rules
Appendix C Case Study
We present two examples involving intricate e-commerce business logic to illustrate the effectiveness of Adaptive Guided Replay. By providing guidance tailored to the model’s needs, it successfully learns to explore and apply starkly different reasoning rules, as demonstrated in Table 7 and Table 8.
Table 7. Impact of Adaptive Guided Replay
Case Study: Impact of Adaptive Guided Replay on a query for “cashmere clothing”. Original Prompt Response Query: cashmere clothing.
@@ 2. Item: This cashmere coat …
4. Attribution Match: … Undyed Cashmere Coat contains cashmereThe conclusion is Excellent.
5. Judgement: … Relevance label is 4-Excellent.
Prompt with Guidance Response Query: cashmere clothing.
@@
4. Attribution Match: … Undyed Cashmere Coat contains cashmere, but the content is below 50%The conclusion is Mismatch.
5. Judgement: … Relevance label is 2-Mismatch.
Table 8. Impact of Adaptive Guided Replay
Case Study: Impact of Adaptive Guided Replay on a query for “chiffon dress”. Original Prompt Response Query: chiffon dress.

### v3

@@ 3.3. Adaptive Guided Replay
While Rule-aware Reward Shaping provides fine-grained feedback, reward sparsity remains a significant challenge, particularly for novel, long-tail queries involving composite rules. This issue is best illustrated by the asymmetry in e-commerce search rules. For instance, a cashmere clothing item is rated 4-Excellent only if its cashmere content exceeds 50%, whereas for chiffon, mere presence suffices. Such composite, asymmetric constraints demand multi-step, fine-grained reasoning, yet models refined via SFT or DPO typically exhibit contracted action spaces and limited exploration. Consequently, when an LLM outputs the correct final label without articulating the intermediate logic, it receives no credit for critical sub-tasks. The resulting learning signal becomes sparse precisely where reasoning is most demanding, thereby degrading credit assignment and hindering convergence on complex e-commerce rules.
@@
To prevent the model from engaging in reward hacking or degrading its reasoning abilities when given complete answers, we adopt a “minimal sufficient guidance” strategy. AGR operationalizes this by performing a fine-grained, on-demand diagnosis. Leveraging the multi-dimensional signals from our rule-aware reward, AGR first calculates the in-batch average reward for each dimension (e.g., category, attribute, and relevance) and compares it against a guidance trigger threshold $\tau$. For any dimension where the reward falls below this threshold, AGR generates targeted guidance $\mathcal{G}(x,y)$ for the selected sparse-reward samples $(x,y)$. This guidance is then formatted into a text prompt and appended to the original input $P(x)$ to form an augmented prompt $P^{\prime}(x)$. The model performs a “guided replay” by regenerating its output based on $P^{\prime}(x)$, allowing it to explore higher-reward reasoning trajectories and fill gaps in its understanding of fine-grained discriminative logic. This dimension-specific intervention effectively mitigates reward sparsity while preserving the model’s autonomous reasoning process and suppressing reward hacking.
4. Experiments
@@ 4.1. Experimental Setup
For offline evaluation, we report per-class F1 for Class-1 (1-Irrelevant), Class-2 (2-Mismatch), Class-3 (3-Related), and Class-4 (4-Excellent), along with Good F1, Macro F1, and Accuracy. Specifically, Good F1 is computed by collapsing Class-3 and Class-4 into a single Good category. This aligns our evaluation protocol with the online business objective, where these high-relevance levels are treated as a unified Good Tier. To validate the real-world impact of our model beyond offline metrics, we conducted online side-by-side human evaluations. This rigorous process is designed to capture user-facing performance differences by directly assessing model outputs on three principal metrics:
  • Good/Same/Bad (GSB): This metric quantifies the relative performance gain of our model via a pairwise comparison. For the same query, human annotators are presented with two Search Engine Result Pages (SERPs) side-by-side—one from our model and one from the baseline—and label our model’s result as Good (superior), Same (equivalent), or Bad (inferior). A GSB score of +xx% signifies that our model produced a superior SERP for xx% more queries compared to the baseline.
  • Query Goodrate: This is a holistic, SERP-level metric assessing the absolute quality of search results. Annotators classify the overall utility of an entire SERP into Good, Mid, or Bad tiers. The final metric is the percentage of queries whose SERPs are judged as either Good or Mid, providing a high-level measure of user satisfaction per query.
@@
All experiments related to TaoSR-AGRL were implemented based on the open-source reinforcement learning framework ROLL(Wang et al., 2025). We used a training batch size of 64, a maximum input length of 2048, and a maximum generation length of 4096. The learning rate was set to 1e-6, with 16 gradient accumulation steps. For each input sample, 16 candidate responses were generated using a sampling temperature of 0.99 and a top-k of 100. For difficulty filtering, an online difficulty threshold range of [0.01, 0.9] was used. To enhance training stability, the PPO clipping range was set to $1\pm 0.2$, value function clipping to 0.5, and Advantage value clipping to $\pm 2.0$. The guidance trigger threshold $\tau$ was 0.1, and the hyperparameter weights $w_{\text{cate}}$, $w_{\text{attr}}$, and $w_{\text{reasoning}}$ were set to 0.4, 0.4, and 0.2 through preliminary experiments to balance different reward components effectively.
@@ 4.4. Ablation Study
We conducted a series of ablation studies on the W-Eval benchmark to systematically analyze the performance contributions of each component in TaoSR-AGRL. The main results are summarized in Table 2. In our analysis, the suffix -RRS denotes the variant equipped with Rule-aware Reward Shaping. The suffix -FG (Fixed Guidance) denotes a special case of our Adaptive Guided Replay where the guidance is fixed to the ground-truth relevance labels.
Synergy of Gating and Reasoning Rewards.
To validate the design of Rule-aware Reward Shaping, we ablated its two pillars: the validity gate and the reasoning reward ($R_{\mathrm{reason}}$). Beyond end-task metrics, we use the Rule Adherence Rate (RAR) to measure if the model’s reasoning is faithful to its own intermediate steps. Ablating the hard validity gate—either by removing or softening it—introduces reward leakage into invalid trajectories. This not only degrades end-task performance but also lowers RAR, thereby confirming that a strict gate is essential for providing a clean, focused training signal. More critically, removing $R_{\mathrm{reason}}$ causes a steep decline in RAR while leaving end-task metrics largely intact. This result provides clear evidence of reward hacking: the model learns to find the correct answer via logically flawed shortcuts. This demonstrates the synergy between our components: the hard gate ensures learning is confined to successful outcomes, while $R_{\mathrm{reason}}$ ensures the reasoning process leading to those outcomes is logically sound and faithful. A detailed breakdown of these ablations is available in Appendix A.
@@
In reinforcement learning, policy entropy is a critical measure of exploration. High entropy is essential for deep reasoning tasks, as it encourages the model to explore a diverse set of inferential strategies. We compare TaoSR-AGRL against TaoSR-AGRL-Static, a variant that applies guidance indiscriminately to all samples. As illustrated in Figure 2(c), TaoSR-AGRL consistently maintains a higher policy entropy. This suggests that excessive or non-adaptive guidance can lead to policy collapse(Juliani and Ash, 2024): the model becomes overly reliant on the guidance, causing its output distribution to become over-confident and collapse into a narrow set of familiar responses. This dependency ultimately stifles the model’s exploratory capabilities and impairs its ability to generalize.
@@
To determine the optimal frequency of intervention, we analyze the model’s overall performance across different values of the replay trigger threshold $\tau$. As detailed in Appendix B, performance peaks at $\tau=0.1$. This result highlights a fundamental trade-off between corrective intervention and policy autonomy. A low $\tau$ provides insufficient correction for challenging instances, hindering learning. Conversely, a high $\tau$ leads to over-reliance on external guidance, risking the policy collapse discussed previously. This over-reliance can impair the model’s intrinsic reasoning ability, even on simpler tasks.
@@ 4.5. Online Evaluation
We validated real-world effectiveness via large-scale online side-by-side human evaluations in production. TaoSR1 served as the baseline bucket and TaoSR-AGRL as the test bucket. We sampled 2,000 authentic user queries and compared the top-10 retrieved results from both buckets in an item-level head-to-head comparison. As shown in Table 3, TaoSR-AGRL consistently outperformed the baseline bucket, achieving average gains of +7.11% in GSB, +1.74 pt in Query Goodrate, and +3.74 pt in Item Goodrate. Improvements are most pronounced for reasoning-intensive traffic (e.g., Q&A and knowledge-based queries), indicating that the proposed model effectively integrates e-commerce domain knowledge with rule-driven reasoning to better resolve complex user intent and specific constraints.
To ensure production readiness for high-throughput serving, we implemented engineering optimizations targeting stable response under load, improved resource utilization, and reduced end-to-end latency. Specifically, we optimized model-level load balancing to mitigate traffic skew, reduced cross-datacenter scheduling overhead to lower tail latency, and applied model quantization to accelerate inference while maintaining ranking quality. With these optimizations, Model FLOPs Utilization (MFU) increased to 35% during training, and the P99 tail latency was reduced to $\leq 400$ ms, enabling reliable large-scale inference.
Despite the relevance gains verified by human evaluation, an initial online deployment revealed a slight degradation in critical business metrics (e.g., orders and GMV). Root-cause analysis traced this issue to the upstream recall stage, which disproportionately retrieved items with strong semantic matching but low sales velocity, thereby reducing conversion efficiency. This observation highlights a common trade-off in e-commerce search: optimizing recall purely for semantic relevance can suppress commercially competitive items whose textual matches are less explicit.
To address this misalignment, we introduced a multi-channel recall mechanism that incorporates personalized signals, along with a pre-ranking stage that jointly optimizes relevance and commercial potential. We then conducted a 4-day online A/B test on 4% traffic over eligible queries. The experiment achieved Direct Clean GMV +0.34%, Direct Clean Orders -0.72%, Direct Clean IPV +3.89%, and Exposed PV +3.86%. Overall, these results show consistent gains in user experience as measured by IPV and PV, with no material degradation of core business metrics. This underscores the necessity of co-designing recall and ranking to balance semantic quality and commercial objectives in end-to-end search systems.
@@ 5. Conclusion
We propose TaoSR-AGRL, an adaptive guided reinforcement learning framework that systematically addresses reward sparsity and exploration stagnation in LLM-based e-commerce search relevance. It integrates Rule-aware Reward Shaping, which transforms sparse terminal outcomes into dense rewards to encourage faithful reasoning, and Adaptive Guided Replay, which provides targeted, on-demand guidance to steer exploration toward high-value trajectories for challenging instances. Extensive offline and online evaluations show that TaoSR-AGRL consistently outperforms strong baselines, improving relevance accuracy and rule adherence while also boosting user engagement in real-world business settings. Deployed on Taobao serving hundreds of millions users, it provides a verified framework for applying reinforcement learning to build robust, industrial-scale reasoning systems.
@@ Appendix A Ablation Study on Rule-aware Reward Shaping
We conducted an ablation study to validate the design of our Rule-aware Reward Shaping. The study evaluates two key components: (1) the validity gate ($\mathbb{I}_{\mathrm{Gate}}$), which selectively applies rewards, and (2) the reasoning reward ($R_{\mathrm{reason}}$), which promotes logical consistency. TaoSR-AGRL employs a hard gate (Gating Coefficient $\lambda=0$), meaning no reward is assigned if the trajectory fails to satisfy the gating condition. To assess whether the model derives relevance in accordance with the operational rules, we report the Rule Adherence Rate (RAR). RAR measures the proportion of predictions where the final Relevance tier is consistent with the predicted Category and Attribute tiers under the Relevance Derivation Rules in Table 6. Consequently, RAR quantifies rule-following behavior and complements standard performance metrics (F1, Accuracy): a model may achieve high Accuracy while violating internal rules, a discrepancy RAR is designed to reveal.
@@
As shown in Table 4, TaoSR-AGRL outperforms all ablated variants and attains the highest RAR. We observe three consistent trends: (1) relaxing or removing the gate causes reward leakage into invalid trajectories, which not only degrades task metrics but also reduces RAR, indicating inferior rule adherence; (2) increasing the softness of the gate exacerbates this leakage and further diminishes internal consistency, underscoring the necessity of a hard gate to concentrate learning on valid reasoning chains; (3) removing $R_{\mathrm{reason}}$ yields notably lower RAR despite similar primary metrics, revealing shortcut solutions—a form of reward hacking—that bypass the intended logic. Collectively, these findings demonstrate that the hard gate provides a strict supervision boundary, while $R_{\mathrm{reason}}$ regularizes the solution space to ensure faithful, rule-consistent derivations, thereby improving both utility and reliability.
@@
Appendix B Impact of the Replay Trigger Threshold
Table 5 presents the sensitivity analysis of the replay trigger threshold $\tau$ on the In-the-Wild evaluation set. The table reports Macro F1 and Accuracy scores as $\tau$ varies within the range of $[0,0.5]$, demonstrating that optimal performance is achieved at $\tau=0.1$.
Table 5. Sensitivity analysis of the Replay Trigger Threshold $\tau$ on the In-the-Wild evaluation set.
Threshold ($\tau$) Macro F1 Accuracy
0 60.23 73.58
0.1 61.53 74.28
0.2 59.91 74.10
0.3 59.40 73.77
0.4 59.45 73.64
0.5 58.70 74.27
Appendix C Relevance Derivation Rules
Table 6 summarizes the Relevance Derivation Rules currently employed on Taobao. Given specific Category and Attribute tiers, the corresponding relevance tier is deterministically mapped as shown.
@@ Appendix C Relevance Derivation Rules
Appendix D Case Study
We present two examples involving intricate e-commerce business logic to illustrate the effectiveness of Adaptive Guided Replay. By providing guidance tailored to the model’s needs, the model successfully learns to explore and apply distinct reasoning rules, as demonstrated in Table 7 and Table 8.
Table 7. Case Study 1: Impact of Adaptive Guided Replay
Case Study: Impact of Adaptive Guided Replay on a query for “cashmere coat”. Original Prompt Response Query: what is cashmere clothing?
@@ 2. Item: This cashmere coat …
4. Attribute Match: … Undyed Cashmere Coat contains cashmereThe conclusion is Excellent.
5. Judgment: … Relevance label is 4-Excellent.
Prompt with Guidance Response Query: what is cashmere clothing?
@@
4. Attribute Match: … Undyed Cashmere Coat contains cashmere, but the content is below 50%The conclusion is Mismatch.
5. Judgment: … Relevance label is 2-Mismatch.
Table 8. Case Study 2: Impact of Adaptive Guided Replay
Case Study: Impact of Adaptive Guided Replay on a query for “chiffon dress”. Original Prompt Response Query: recommend a chiffon dress suitable for a seaside vacation.